Fermi National Laboratory
 

Accelerator Update for November 15, 2006 - November 17, 2006

 

Wednesday November 15

The day shift began with Operations monitoring the Tevatron (TeV) store 5066 (36.92E30) and a Recycler stash (42.3E10), but with no stashing due to a Recycler damper problem, with the Antiproton Source (Pbar) stacking (142.7E10), with Linac, Booster, and the Main Injector (MI) operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the store, and with MiniBooNE taking beam. NuMI was off due to chiller repairs.

Operations prepared the Recycler and Pbar at 11:22 AM to stash antiprotons. Experts conducted studies during this transfer.

Operations terminated store 5066 at 12:06 PM.

The Recycler stash was completed at 12:35 PM.

Operations prepared CDF and D0 for accesses at 12:39 PM.

The Recycler began another stash at 2:55 PM.

At 3:36 PM, Operations reported that the stash had ended, but Pbar had a problem resuming stacking. The lithium lens had tripped off three times. Operations contacted experts.

At 4:33 PM, a Linac RF station (LRF5) was off due to a permanent inhibit. Experts reported that the LRF5 power amplifier (PA) was sparking.

Linac experts reported at 5:33 PM that LRF5 was up and running, but their investigations were continuing.

Operations put the TeV into shot setup at 6:03 PM.

At 7:41 PM, shortly after the antiprotons had been loaded into the TeV, the TeV quenched at sector F3. The Tevatron Electron Lens (TEL2) had tripped off just before the quench. After a lot of investigation, experts thought the cause of the quench was due to an orbit drift at F32.

Operations put the TeV through a dry squeeze at 11:05 PM.

The Recycler completed a stash at 11:51 PM.

Thursday November 16

The midnight shift began with Operations monitoring a stash (73.8E10), with Pbar stacking (18.6E10), and with MiniBooNE taking for beam.

Operations loaded the TeV with a 14x0 (14 bunches of protons and no antiprotons) study store at 12:26 AM.

Operations terminated the 14x0 study store at 1:39 AM. A TeV expert began a new study.

At 4:54 AM, a serious problem with Pbar beam position occurred at AP50.

The Recycler completed a stash at 4:57 AM.

At 5:07 AM, the AP50 beam position problem got worse and continued for over an hour and a half. Operations contacted experts.

At 6:32 AM, Operations reported that the problem at AP50 was due to sparking in a Debuncher kicker (D:IKIK). An expert investigated and said the repair would take about two hours. Experts suggested that it would be better to stash the antiprotons before beginning the repairs.

The Recycler completed a stash at 7:38 AM.

Pbar experts reported at 8:41 AM that the kicker was fixed. Pbar resumed stacking.

Operations loaded 14x0 study store into the TeV at 10:11 AM. TeV experts used the store to measure and optimize orbits and check the beam position F32.

Operations loaded another 14x0 study store at 11:16 AM.

At 11:54 AM, Operations reported a BPM problem in the TeV’s F sector. Experts investigated and discovered a bad control card in a VME crate.

The Recycler completed a stash around noon.

Pbar experts began a stacktail study at 1:35 PM.

Operations put the TeV into shot setup at 1:43 PM.

At 2:24 PM, the Linac 55 LCW system had a pressure and temperature drop that caused LRF3 to trip. A technician at CUB said a mixing control loop had hung up.

Linac resumed sending beam at 2:49 PM even though the 55 LCW system was still unstable.

Operations established store 5074 at 4:29 PM with an initial luminosity of 132.5E30.

NuMI experts reported at 5:05 PM that a bad contactor switch would keep them from running tonight.

At 6:32 PM, a TeV separator spark caused the luminosity to dip slightly.

At 7:10 PM, Operations reported a power glitch. The stack and stash survived, but the TeV quenched at multiple houses. The store was lost. Most of Linac and Booster tripped off. Feeder 47 was the problem. Operations contacted the High Voltage Coordinator.

The High Voltage Coordinator reported at 8:55 PM that the 38/39 Well Pump #3 went down and caused the glitch on feeder 47. They will disconnect the pump and then switch the power loads to other feeders.

Operations prepared CDF and D0 for an access into their collision halls at 9:33 PM.

The Booster RF system was brought up at 9:52 PM.

At 10:23 PM, Operations reported that all the TeV pond and LCW pumps had tripped off during the glitch.

A Linac expert had all the Klystron RF stations on by 11:30 PM.

Friday November 17

The midnight shift began with Operations monitoring a stash (58.8E10) and stack (26.7E10), with NuMI off for switch repairs, with MiniBooNE off due to DAQ problems, and with all other machines recovering from the glitch.

A TeV RF expert reported at 12:11 AM that the TeV RF system was ready.

D0 and CDF were out of access by 12:25 AM.

Operations began the TeV checkout at 1:19 AM.

The Linac beam came on at 1:38 AM. An expert had to switch to the H- Source due to problems with the I- Source.

Operations established beam in the Booster at 1:42 AM.

At 1:51 AM, MiniBooNE experts reported that they were having problems with data acquisition. They cannot take beam until this problem gets fixed.

Operations ramped the TeV to 980 GeV at 3:44 AM.

Operations put the Recycler into shot setup to stash antiprotons at 4 AM.

At 5:54 AM, Operations reported that only two of the five transfers of antiprotons made it into the Recycler. Later, experts discovered that a power supply (I:HV703) was not regulating correctly. The Recycler must have this power supply to stash.

At 6:05 AM, a TeV expert reported that some of the low beta magnet leads were very cold. The low beta power supplies were being brought up slowly to warm the leads.

At 7:55 AM, D0 experimenters reported that they needed to access their hall again. Operations prepared D0 for access.

The Plans for Friday and the Weekend

The plans for today are to allow a TeV transfer Hall access for repairs and for all other machines to complete their recovery from the glitch. Operations hopes to establish a store in the TeV tonight. The weekend will consist of stacking, stashing, and storing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/17/2006    email Fermilab